# Persistent Autoregressive Mapping with Traffic Rules for Autonomous Driving

### v1

@@ Abstract
Preprint. Under review.
Introduction
@@
Related Work
HD Map Construction
HD map construction has evolved significantly towards end-to-end solutions (Li et al. 2024; Zhang et al. 2024, 2023). Large-scale autonomous driving datasets (Wilson et al. 2021; Caesar et al. 2020) and specialized traffic sign datasets (Behrendt, Novak, and Botros 2017; Stallkamp et al. 2012; Yu et al. 2020; Fregin et al. 2018; Zhu et al. 2016) have played a crucial role in advancing this field. Early works like VectorMapNet (Liu et al. 2023) pioneered sequential vector prediction through a two-stage approach: first predicting key points as a set prediction task, then generating map elements sequentially. MapTR (Liao et al. 2023a) further advanced this field by introducing a permutation-equivalent modeling approach, representing map elements as point sets with equivalent permutations to avoid shape ambiguity. However, these approaches focus solely on geometric elements, remaining “semantically blind” to traffic rules.
Recent attempts to incorporate traffic rules have emerged, but with significant limitations. OpenLane-V2 (Wang et al. 2023) treats rule assignment as simple classification, while MapDR (Chang et al. 2025) requires pre-constructed vectors, preventing true end-to-end integration. Meanwhile, newer architectures like PivotNet (Ding et al. 2023) have explored unified pivot-based representations and point-to-line mask modules to better capture geometric relationships. A critical challenge remains: modeling the persistent effectiveness of traffic rules beyond their visible range while maintaining geometric-semantic consistency. This gap motivates our work towards unified geometric-semantic mapping.
MLLMs in Driving
The sequential nature of driving requires continuous interpretation of both geometric and semantic information. Traditional approaches often fragment this sequential process, leading to disconnected representations. While recent works have explored sequential modeling through various architectures - from VectorMapNet’s autoregressive decoder (Liu et al. 2023) to MapTR’s unified Transformer structure (Liao et al. 2023a) - few address the temporal persistence of semantic rules. The evolution from two-stage approaches to end-to-end frameworks highlights the field’s progression towards more efficient and integrated solutions.
MLLMs have shown promise in sequential reasoning across various domains (Anthropic 2024; Bai et al. 2023; OpenAI 2024). Their success in understanding spatial relationships and maintaining contextual awareness suggests potential for HD mapping (Cui et al. 2024; Xu et al. 2024; Sima et al. 2024; Tian et al. 2024; Chen et al. 2024b; Ding et al. 2024; Choudhary et al. 2024; Cao, Wei, and Ma 2025; Zeng et al. 2025, 2024; Xie et al. 2025; Wei et al. 2024). Particularly relevant is their ability to ”narrate” sequential information while maintaining long-term dependencies, a crucial capability for persistent rule-aware mapping. This aligns well with the current trend towards unified, efficient architectures in HD map construction, suggesting a promising direction for integrating geometric and semantic understanding in autonomous driving.
Method
@@
Vectorized Lane Accuracy. To assess predicted lane vectors $\hat{V}$ against ground-truth $V$, we implement topology-aware Intersection-over-Union (IoU  (Zheng et al. 2020)) evaluation. For each lane polyline $\hat{l}_{i}\in\hat{V}$, we rasterize lanes into fixed-wide (6m in our experiments) binary masks through polyline expansion. The IoU between each predicted mask $\hat{l}_{i}$ and ground-truth masks $\left\{l_{i}\right\}_{i=1}^{k}$ is computed, with matches established when IoU >> 0.5. We define the vector accuracy metric $\mathcal{F}_{\text{vec}}$ as Eq 1:
@@
where $f$ denotes the MLLM, and $Prompt^{*}$ represents optional prompts that enable both continuous map-rule construction ($P_{con}$ in Sec Map-Rule Cache) and interactive ($P_{rule}$ in Sec Interactive Prompt). These components will be thoroughly discussed in the following sections.
Map-Rule Co-Construction
@@
Implementation Details
Dataset and Metric. We evaluate PAMR on MapDRv2, with metrics defined in Sec MapDRv2. The dataset encompasses diverse scenarios, weather conditions, and traffic situations, comprising over 1010k traffic scene segments, 1818k driving rules, and 400400k images with $1960\times 1240$ resolution. Each segment is set to $224\times 224$ ($W\times H$, corresponding to a $22.4m\times 22.4m$ real-world area), with an overlap ratio $\delta$ of $10\%$. The dataset is split into train and test sets with a $9:1$ ratio for evaluation.
@@
Visualization
Our framework demonstrates robust performance across various challenging scenarios, as illustrated in Fig 5. The map-rule mechanism effectively addresses several key challenges in HD map construction. First, it maintains consistency during lane-changing scenarios and generates smooth, accurate curved lane vectors, demonstrating precise vector modeling capabilities. Second, our cache mechanism successfully preserves rule awareness even when traffic signs move beyond the current map-rule. Additionally, the model exhibits strong semantic understanding by accurately associating multiple rules from individual traffic signs with their corresponding lanes. Please refer to the supplementary materials for additional visualizations.
Conclusion

### v2

@@ Abstract
Codehttps://miv-xjtu.github.io/PAMR/
1 Introduction
@@
2 Related Work
2.1 HD Map Construction
Fueled by large-scale autonomous driving (Wilson et al. 2021; Caesar et al. 2020) and traffic sign datasets (Behrendt et al. 2017; Stallkamp et al. 2012; Yu et al. 2020; Fregin et al. 2018; Zhu et al. 2016), HD map construction is rapidly evolving towards end-to-end solutions (Li et al. 2024; Zhang et al. 2024, 2023; Yuan et al. 2025). Pioneering works like VectorMapNet (Liu et al. 2023) and MapTR (Liao et al. 2023a) established methods for sequential and permutation-equivalent vector prediction. However, despite architectural advances like PivotNet (Ding et al. 2023) for improving geometry, these methods remain “semantically blind” to traffic rules.
Recent attempts to incorporate semantics are limited. OpenLane-V2 (Wang et al. 2023) treats rule assignment as simple classification, while MapDR (Chang et al. 2025) relies on pre-constructed vectors, breaking the end-to-end paradigm. A critical challenge therefore remains: modeling the persistent effectiveness of traffic rules beyond their visible range while ensuring geometric-semantic consistency. This gap motivates our work towards unified geometric-semantic mapping.
2.2 MLLMs in Driving
Driving’s sequential nature requires continuous interpretation of semantics, yet even advanced sequential models like VectorMapNet (Liu et al. 2023) and MapTR (Liao et al. 2023a) largely neglect the temporal persistence of traffic rules. This is where Multimodal Large Language Models (MLLMs) show great promise, known for their success in sequential reasoning and contextual awareness (Anthropic 2024; Bai et al. 2023; OpenAI 2024).
Their potential for understanding complex driving scenarios is being actively explored (drivellm; Xu et al. 2024; Sima et al. 2024; Tian et al. 2024; Chen et al. 2024b; Ding et al. 2024; Choudhary et al. 2024; Cao et al. 2025; Zeng et al. 2025a, 2024; Xie et al. 2025b; Wei et al. 2024; Xie et al. 2025a; Zeng et al. 2024, 2025a, 2025b). Particularly relevant is their ability to “narrate” sequential information while maintaining long-term dependencies—a crucial capability for the persistent rule-aware mapping that current methods lack. This aligns with the trend towards unified architectures, offering a promising path to integrate deep geometric and semantic understanding in autonomous driving.
3 Method
@@ 3.1 MapDRv2
Vectorized Lane Accuracy. To assess predicted lane vectors $\hat{V}$ against ground-truth $V$, we implement topology-aware Intersection-over-Union (IoU  (Zheng et al. 2020)) evaluation. For each lane polyline $\hat{l}_{i}\in\hat{V}$, we rasterize lanes into fixed-wide binary masks through polyline expansion. The IoU between each predicted mask $\hat{l}_{i}$ and ground-truth masks $\left\{l_{i}\right\}_{i=1}^{k}$ is computed, with matches established when IoU >> 0.5. We define the vector accuracy metric $\mathcal{F}_{\text{vec}}$ as Eq 1:
@@ 3.2 PAMR
where $f$ denotes the MLLM, and $Prompt^{*}$ represents optional prompts that enable both continuous map-rule construction ($P_{con}$ in Sec 3.4) and interactive ($P_{rule}$ in Sec 3.5). These components will be thoroughly discussed in the following sections.
3.3 Map-Rule Co-Construction
@@ 4 Experiments
4.1 Implementation Details
Dataset and Metric. We evaluate PAMR on MapDRv2, with metrics defined in Sec 3.1. The dataset encompasses diverse scenarios, weather conditions, and traffic situations, comprising over 1010k traffic scene segments, 1818k driving rules, and 400400k images with $1960\times 1240$ resolution. Each segment is set to $224\times 224$ ($W\times H$, corresponding to a $22.4m\times 22.4m$ real-world area), with an overlap ratio $\delta$ of $10\%$. The dataset is split into train and test sets with a $9:1$ ratio for evaluation.
@@ 4.3 Ablation Study
4.4 Visualization
Our framework demonstrates robust performance across various challenging scenarios, as illustrated in Fig 5. The map-rule mechanism effectively addresses several key challenges in HD map construction. First, it maintains consistency during lane-changing scenarios and generates smooth, accurate curved lane vectors, demonstrating precise vector modeling capabilities. Second, our cache mechanism successfully preserves rule awareness even when traffic signs move beyond the current map-rule. Additionally, the model exhibits strong semantic understanding by accurately associating multiple rules from individual traffic signs with corresponding lanes.
5 Conclusion
@@ 6 Limitations
Acknowledgments
This work was support by the National Natural Science Foundation of China No. 62572385, the Fundamental Research Funds for the Central Universities No. xxj032023020, and CAAI-CANN Open Fund, developed on OpenI Community.